the average, or mean D of three exam grades x, m and n is given by the formula
D= x+m+n
3
(a) solve the formula for n
(b) use the formula in part (a) to solve the problem.
On your first two exams, your grades are 84% and 87%. x= 84 and m=87%. what must you get on the third exam to have an average of 91%?
please explain how to determine the answer

Answer by stanbon(75887)

the average, or mean D of three exam grades x, m and n is given by the formula
D = (x+m+n)/3
---
(a) solve the formula for n
x+m+n = 3D
n = 3D-x-m
------------------------------
(b) use the formula in part (a) to solve the problem.
On your first two exams, your grades are 84% and 87%. x= 84 and m=87%. what must you get on the third exam to have an average of 91%?
please explain how to determine the answer
---
Let the average on the 3rd exam be "x".
---
Equation:
(0.84+0.87+x)/3 = 0.91
1.71 + x = 3*0.91
x = 2.73-1.71
x = 1.02
---
You would have to average 102% on the final exam to have
an average on 91% on the three exams.
-----------------
Cheers,
Stan H.